how many grams of NaF are produced from 3.8 moles of Na?


1
Expert's answer
2021-03-10T06:40:31-0500

Q170127

how many grams of NaF are produced from 3.8 moles of Na?


Solution:

Step 1:

Let us write the reaction first.

Na(s) reacts with F2 (g) to give NaF (s)


Na(s) + F2 (g) ⟹ NaF (s)


balance this reaction first.

There is 2 F on left side and 1 F on right side.

balance the F by writing proper coefficient in front of NaF.


Na(s) + F2 (g) ⟹ 2 NaF (s)


Now we can see that there is 1 Na on left side and 2 Na on right side.

Balance the Na by writing proper coefficient in front of Na.


2 Na(s) + F2 (g) ⟹ 2 NaF (s)


Now the reaction is balanced for all the elements.


Step 2 : Find the moles of NaF that can be formed from 3.8 moles of Na.


The mole to mole ratio of Na and NaF in the balanced reaction is 2 : 2 = 1 : 1


so moles of NaF formed from 3.8 moles of Na is

= "3.8 \space moles \space of \space Na * \frac{1 \space mol \space NaF }{1 \space mol \space Na } = 3.8 \space moles \space of \space NaF."


Step 3 : To convert 3.8 moles of NaF to grams by using molar mass of NaF.


molar mass of NaF = 1 * atomic mass of Na + 1 * atomic mass of F

= 1 * 22.99 g/mol + 1 * 18.998 g/mol

= 22.99g/mol + 18.998 g/mol

= 41.988 g/mol


grams of NaF = "3.8 \space moles \space of \space NaF * \frac{41.988 \space g \space of \space NaF }{1 \space mol \space of \space NaF } = 159.55grams."


In the question, we are given 3.8 moles of Na in 3 significant figure, so our final answer must also be in 3 significant figures.


Hence the mass of NaF produced from given moles of Na is 160 grams.


Thank you. :)